OpenStudy (anonymous):

Please help!!!!! the distance a body falls varies directly as the square of the time that it falls. If it falls 867 meters in 17 seconds how far will it fall in 4 seconds? Please show me how you did this! I need to know how to do it Thank you!

OpenStudy (anonymous):

Hi, I think it is as follows: If d= distance(m), and k= constant, and t = time (sec) Using the info above, it is true that d=kt^2 (where ^2 means squared) Initially, using the figures above, 867 = k x 17^2 = k x 289 Therefore k = 3. So if t = 4, you have d = 3 x 4^2 = 3 x 16 = 48m. Hope this helps :D

OpenStudy (anonymous):

Oh thank you so very much!!!! :) I have one more problem if you can help

OpenStudy (anonymous):

yea sure post it !! :D

OpenStudy (anonymous):

as a new question!!

OpenStudy (anonymous):

ok yellows very inversely as grays cubed. When there are 270 yellows there are 29 grays how many yellows are present when there are 6 grays?

OpenStudy (anonymous):

ok